### v1

# How precisely can we measure the ages of subgiant and giant stars?

Abstract
Precise stellar ages are fundamental to Galactic archaeology. However, obtaining reliable age estimates and uncertainties for field stars has been a long-standing challenge. We test the fidelity of ages from recent catalogs of giants and subgiants using wide binaries, whose components formed at the same time and thus should have consistent inferred ages. We find that subgiant ages based on spectroscopic metallicities from Xiang & Rix (2022) are generally consistent within their reported uncertainties, implying that fractional uncertainties of are realistically achievable. In contrast, we find that published photometric subgiant ages and red giant ages underestimate true uncertainties by factors of . These results demonstrate that accurate metallicity and -element abundances are essential for precise isochrone ages and establish wide binaries as a powerful, model-independent benchmark for calibrating stellar age measurements in the era of large spectroscopic surveys.
1 Introduction
Reconstructing the Milky Way’s formation and chemical enrichment history requires precise ages for large numbers of stars across the Galaxy (e.g., Twarog, 1980; Soderblom, 2010; Casagrande et al., 2016; Sanders & Das, 2018; Deason & Belokurov, 2024). Modern surveys now routinely estimate ages for millions of field stars by combining astrometry, multi-band photometry, spectroscopy, and, where available, asteroseismology (e.g., Xiang et al., 2017; Queiroz et al., 2018; Sanders & Das, 2018; Helmi, 2020; Xiang & Rix, 2022; Nataf et al., 2024; Wang et al., 2023, 2025). Among different evolutionary stages, subgiants are especially powerful chronometers: their luminosities depend sensitively on core mass, which scales with stellar mass and thus main-sequence lifetime. With sufficiently precise parallaxes and metallicity measurements, isochrone fitting can theoretically achieve relative age precisions at the percent level for subgiant populations (e.g., Nataf et al., 2024).
For Galactic archaeology, achieving stellar ages with precision over a wide age range ( Gyr) is critical for reconstructing the Galaxy’s star formation history and quantifying its chemical evolution (e.g., Xiang et al., 2017; Bonaca et al., 2020; Helmi, 2020). However, estimating accurate age uncertainties remains challenging. Even small differences in how ages are estimated can lead to different conclusions about the timeline of the Galaxy’s disk and halo assembly (e.g., Conroy et al., 2022; Xiang & Rix, 2022). Many internal validation tests rely on stellar evolution models, often the same ones used to derive the ages, which limits their diagnostic power and fails to probe systematic errors fully. Empirical checks using open and globular clusters provide an important comparison, but such clusters span a narrow range of ages and metallicities. An independent benchmark that samples a broader range of stellar populations is therefore essential.
Wide binaries provide such a benchmark. They are distributed across the Galactic disk and halo, and effectively sample the entire range of stellar ages and metallicities found in the Milky Way. Their components share a common chemical composition (e.g., Hawkins et al., 2020) and are coeval (i.e., effectively the same age; Makarov et al., 2008; Kraus & Hillenbrand, 2009). This makes them uniquely valuable for testing whether reported age uncertainties are realistic: discrepancies between the two components, normalized by the quoted errors, directly reveal whether catalog error bars are under- or overestimated. The coeval and co-chemical nature of wide binaries has enabled progress across many areas of stellar astrophysics, including the calibration of gyrochronology (e.g., Barnes, 2007; Mamajek & Hillenbrand, 2008; Deacon et al., 2016; Otani et al., 2022; Gruner et al., 2023; Silva-Beyer et al., 2023; Lares-Martiz et al., 2024), M-dwarf metallicity estimates (e.g., Bonfils et al., 2005; Lépine & Bongiorno, 2007; Johnson & Apps, 2009; Rojas-Ayala et al., 2010; Mann et al., 2013; Montes et al., 2018), stellar activity-age relations (e.g., Garcés et al., 2011; Chanamé & Ramírez, 2012), consistency tests for chemical tagging (e.g., Andrews et al., 2019; Hawkins et al., 2020; El-Badry et al., 2021), the age–metallicity relation (Rebassa-Mansergas et al., 2016), and the initial–final mass relation for white dwarfs (Zhao et al., 2012; Andrews et al., 2015; Hollands et al., 2024). For a recent review discussing wide binaries with Gaia, see El-Badry (2024).
In this work, we use wide binaries in which both components are evolved stars with independently inferred ages to test the accuracy of reported age uncertainties in a model-independent way. The remainder of this paper is organized as follows. In Section 2, we describe the sample selection methodology. Section 3 presents our main results, and Section 4 provides a discussion of our results. Finally, Section 5 summarizes our conclusions and considers future directions.
2 Data and Sample Selection
2.1 Parent catalogs
Catalogs providing ages for a large sample of evolved stars have recently been constructed using both photometric and spectroscopic data. The three catalogs that we focus on are those of Xiang & Rix (2022), Nataf et al. (2024), and Wang et al. (2023). We briefly summarize each below.
Xiang & Rix (2022) constructed a subgiant catalog by combining Gaia astrometry with LAMOST DR7 spectroscopy. They first select subgiants using an HR diagram cut. Stellar atmospheric parameters (, [Fe/H], [/Fe]) are derived for all spectra with SNR , restricting to stars with K, where the spectral fits are more robust. They emphasize that accurate abundances are crucial for dating subgiants, showing that even modest [/Fe] offsets (e.g., dex) can shift isochrone ages by Gyr. Stars with mag were excluded to avoid contamination from He-burning horizontal branch stars. Also, objects whose spectro-photometric and Gaia parallax distances disagreed by more than were removed, as were unresolved binaries flagged by Gaia diagnostics. The final catalog thus minimizes contamination from misclassified evolutionary states and unresolved companions. Ages were inferred by fitting the derived parameters (, , [Fe/H], [/Fe]) together with Gaia+2MASS photometry to a grid of Yonsei-Yale (Y2) isochrones (Yi et al., 2001; Demarque et al., 2004).
Nataf et al. (2024) developed a photometric pipeline to infer subgiant ages from Gaia DR3 distances and UV-IR photometry. Subgiants were selected via a cut in the color-magnitude diagram (CMD), but since their locus overlaps with both the main-sequence turn off (MSTO) and base of the red giant branch (RGB), the cut is strongly age- and metallicity-dependent. To mitigate this, they split the sample into two: a “Primary Sample” targeting , where subgiants are cleanly separated from metal-rich MSTO stars, and a metal-poor “annex” selected using GALEX NUV or SDSS/Skymapper -band photometry. Additional astrometric and photometric quality cuts were applied to enhance purity and remove unresolved binaries. Ages were inferred using the isochrones package (Morton, 2015), which fits MIST isochrones to astrometry, photometry, and (implicitly constrained) metallicity. Unlike Xiang & Rix (2022), metallicities here are not derived spectroscopically but are instead primarily inferred from UV photometry. By comparing to Xiang & Rix (2022), Nataf et al. (2024) find systematic offsets: their metallicities were typically higher by dex, and ages were lower, with an average ratio of . They concluded that their UV photometry offers a comparable diagnostic to large spectroscopic surveys for subgiant age inference. Overall, they report a median age precision in their sample of
Wang et al. (2023) derived ages for LAMOST DR8 RGB and red clump (RC) stars. Giants were selected by K and . To distinguish RGB from RC stars, they employed pseudo-asteroseismic diagnostics (, ) inferred from spectra via a neural network trained on Kepler giant stars with asteroseismology. These quantities, when measured precisely, can effectively distinguish RGBs from RCs (e.g., Bedding et al., 2011; Stello et al., 2013; Pinsonneault et al., 2014; Vrard et al., 2016; Elsworth et al., 2017; Wu et al., 2019). Because such labels exist for many Kepler giants that are also observed by LAMOST, Wang et al. (2023) trained a neural network to predict and directly from LAMOST spectra (following Ting et al. 2018; Wu et al. 2019). This enabled RGB/RC classification and subsequent mass/age determination. For the training set (LAMOST+Kepler), they fit PARSEC isochrones (Bressan et al., 2012) in a Bayesian framework, then applied the trained model to all DR8 giants. Typical uncertainties were (age) and (mass) for RGBs and (age) and (mass) for RCs. Comparisons against open clusters and other catalogs suggested that for spectra with SNR, age uncertainties are for RGBs and for RCs, with mass uncertainties below .
2.2 wide binary cross-match
We assess the reported age uncertainties using Gaia wide binaries. El-Badry et al. (2021) constructed a catalog of 1 million high-confidence wide binaries within kpc of the Sun using Gaia DR3 astrometry. We extend their sample to kpc using the exact same methodology, which preserves the original emphasis on purity while adding another new wide binaries. Because giant stars are intrinsically bright and thus maintain precise astrometry at larger distances, this extension substantially boosts the number of usable systems for this study. The resulting catalog of 1.6 million pairs within kpc serves as the foundation for our analysis, from which we identify binaries where both components are evolved stars and have independently inferred ages from the parent catalogs.
We positionally cross-match each of the parent catalogs to the wide binary sample using a tolerance. From these matches, we select systems in which both components of a wide binary are present in the same catalog and have reported ages. This yields , , and binaries for Xiang & Rix (2022), Nataf et al. (2024), and Wang et al. (2023), respectively. We further subdivide the Nataf et al. (2024) sample to include only stars in their Primary Sample defined above (), and the Wang et al. (2023) sample to those with spectra of SNR (). Also, while not the focus of this study, requiring that only one wide binary component has an age estimate provides , , and systems, which may be useful for other work. The full wide binary table is available in Appendix A.
3 Results
To calibrate uncertainties, we compute the uncertainty-normalized age difference
for each of the samples, and determine its standard deviation, . If quoted errors are realistic and independent, should be distributed as a Gaussian with unit variance and mean . For Nataf et al. (2024), who report asymmetric errors (e.g., 16th/84th percentiles), we symmetrize the uncertainty used in via the RMS of upper and lower errors, . This approach avoids directional bias and provides a single representative for computing , while we retain the asymmetric error bars in visualizations111Omitting this symmetrization does not impact the broader results..
Figure 1 compares the component ages reported in the three catalogs. The top panels show the Kiel diagrams for each sample, plotting both the primary (brighter) and secondary (fainter) stars in each wide binary. The bottom panels compare the ages of the primary (Age 1) and secondary (Age 2) components, with the published uncertainties shown. For the Nataf et al. (2024) catalog, we distinguish between their full sample and the “Primary Sample”, which is restricted to (e.g., top middle of Figure 1). For the Wang et al. (2023) sample, we additionally split their giants by spectral SNR, highlighting the SNR subsample (for both stars), where they report typical age uncertainties of . For each case, we compute and report , the standard deviation of the normalized age difference, and overplot the 1:1 relation for reference.
We find that the catalog of Xiang & Rix (2022) produces the most consistent subgiant ages. For the full sample, we find , which is mostly attributed to a single strong outlier with Age 1 = Gyr and Age 2 = Gyr. Removing this outlier reduces the spread to . The outlier also shows a discrepant [Fe/H] between the two stars, which explains the discrepant age estimates222We check that this source is indeed confidently a wide binary, with a chance alignment probability of (El-Badry et al., 2021) at a physical separation of au and distance pc (Appendix A).. After removing this outlier, the ages of wide binary components are statistically consistent within their reported errors. While this is the smallest wide binary sample of the three (), it is the only one for which .
By contrast, both Nataf et al. (2024) and Wang et al. (2023) appear to underestimate their age uncertainties. For Nataf et al. (2024), the full sample shows substantial excess scatter, and even in the Primary Sample, the dispersion remains well above unity (). This suggests that residual systematics – such as photometric blends, biased extinction values, or metallicity misestimation – inflate the true uncertainties beyond the reported statistical errors. For Wang et al. (2023), the discrepancy is larger still: the full giant sample yields , implying that the uncertainties are underestimated by a factor of 5 on average. Unlike in the Xiang & Rix (2022) sample, it is not possible to substantially reduce the scatter here by removing any one outlier. Restricting to SNR spectra somewhat improves the age precision (), but this still implies that the reported uncertainties are underestimated by a factor of .
In summary, for the Nataf et al. (2024) and Wang et al. (2023) samples indicates that their reported age uncertainties are systematically underestimated. The reduction in scatter from the full Nataf et al. (2024) sample to their Primary Sample, and the improved agreement for high-SNR () sources in Wang et al. (2023), show that stringent quality cuts enhance reliability. Nevertheless, even after these cuts, the true uncertainties remain underestimated by factors of . Among the three age catalogs considered, only Xiang & Rix (2022) achieve age estimates that are both precise and empirically validated by wide binaries. Their success likely stems from the use of spectroscopic [Fe/H] and [/Fe] measurements to anchor the isochrone grids for each star. The median fractional age uncertainty in their sample is .
4 Discussion
4.1 Interpreting wide binary Constraints
The uncertainties inferred from our wide binary analysis represent a lower limit on the true total age uncertainties. Since both components share nearly identical metallicities and similar effective temperatures (when both are evolved), systematics that depend on these parameters are largely canceled out. Likewise, any global offset in the age scale (e.g., all ages being systematically over- or underestimated) would not manifest in our relative test.
These limitations imply that while wide binaries provide a useful test of reported uncertainties, they do not fully capture absolute accuracy. Systematic uncertainties associated with stellar evolution models set another floor for absolute age uncertainties (e.g. Ying et al., 2025), which are crucial for cosmological and chemo-dynamical applications where absolute ages matter, such as in calibrating the age of the Universe or the early enrichment timeline. Our results therefore complement, rather than replace, absolute calibration efforts.
4.2 Ages of Red Giants from [C/N] ratios
Another method of determining stellar ages for red giants relies on chemical clocks: elemental abundance ratios that trace stellar evolution or Galactic chemical enrichment. For red giants, the most widely used method for aging them is the surface carbon-to-nitrogen ratio ([C/N]), which changes during the first dredge-up and subsequent mixing on the RGB. This ratio correlates with stellar mass and therefore with age (e.g., Masseron & Gilmore, 2015; Martig et al., 2016; Roberts et al., 2024). In contrast to abundance ratios such as [Fe/H] or [/Fe], which depend on Galactic chemical-evolution modeling and can introduce population-dependent biases (e.g., Pagel, 2009; Nissen, 2015), [C/N] is directly linked to stellar structure and internal mixing physics.
However, theoretical modeling of mixing and extra-mixing processes remains uncertain, particularly at low metallicities, where these effects can alter the [C/N]-mass relation (e.g., Shetrone et al., 2019; Roberts et al., 2024). Empirical calibration to asteroseismic masses and ages, for example, has proven valuable for anchoring this relation (e.g., Bellinger et al., 2016; Ness et al., 2016; Mackereth et al., 2019; Hon et al., 2020; Anders et al., 2023; Leung et al., 2023; Stone-Martinez et al., 2024, 2025). Currently, these estimates provide reported age precisions of (e.g., Anders et al., 2023; Stone-Martinez et al., 2025).
When cross-matched with our wide binary catalog, we find no binaries in which both components are present in the Anders et al. (2023) APOGEE sample, while seven such pairs exist in the Stone-Martinez et al. (2025) SDSS-V DR19 catalog. For these systems, the inferred age differences are in fact consistent with the reported uncertainties (), considering the small sample size, but the reported uncertainties are quite large, with a median value of . Thus, while the reported errors appear realistic, their true precision is comparable to that of isochrone ages for red giants (e.g., Wang et al., 2023). The continued expansion of spectroscopic surveys, combined with large samples of asteroseismic masses and wide binaries (for calibrating uncertainties), offers an opportunity to refine these abundance-based age relations across metallicity and evolutionary phase.
5 Conclusions
Precise stellar ages are crucial for reconstructing the Milky Way’s star formation, assembly, and chemical-enrichment history, as well as for testing stellar evolution models. As such, several recent studies estimated ages for millions of evolved stars throughout the Galaxy, reporting uncertainties. In this study, we use wide binaries as an empirical benchmark to test the reported age uncertainties of evolved stars. Because the two members of a wide binary are coeval, the average difference between their reported ages directly measures the true uncertainties on such estimates.
Our analysis shows that the subgiant ages of Xiang & Rix (2022) are consistent with their quoted precisions: after removal of a single strong outlier, the scatter between binary components is fully consistent with the reported errors (). In contrast, both the catalogs of Nataf et al. (2024) and Wang et al. (2023) underestimate age uncertainties, with wide binary tests indicating effective errors larger than reported formal errors. Even among their quality-controlled subsamples, we find , demonstrating that additional systematics remain unaccounted for.
A key difference between the subgiant samples of Xiang & Rix (2022) and Nataf et al. (2024) lies in their treatment of metallicity. The former use spectroscopic measurements of both [Fe/H] and (critically) [/Fe], highlighting the importance of accurate chemical abundances for reliable isochrone ages.
This study underscores the utility of wide binaries as a model-independent benchmark for stellar age measurements. Unlike internal validation tests or calibrations to clusters, wide binaries provide a large set of validation experiments across the Galaxy, and most importantly, across a wide range of ages, metallicities, and -abundances, enabling robust empirical assessments of true uncertainties.
Gaia XP spectra provide metallicities for hundreds of millions of stars (e.g., Andrae et al., 2023), which, when combined with effective temperatures and luminosities, may allow ages to be estimated for subgiants without the need for higher-resolution spectroscopy. Extending such XP-based analyses and validating them with wide binaries represents a promising avenue for providing isochrone ages for a larger sample of subgiants. Looking ahead, as surveys deliver larger samples of evolved stars with precise chemical abundances, wide binaries will remain a stringent benchmark for calibrating the precision of stellar age and abundance measurements.
6 Acknowledgments
C.S. acknowledges support from the Department of Energy Computational Science Graduate Fellowship. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, under Award Number DE-SC0026073. This research was supported by NSF grant AST-2307232 and by Scialog grant #SA-LSST-2024-084a from the Research Corporation for Science Advancement and the Heising-Simons Foundation.
This work presents results from the European Space Agency (ESA) space mission Gaia. Gaia data are being processed by the Gaia Data Processing and Analysis Consortium (DPAC). Funding for the DPAC is provided by national institutions, in particular the institutions participating in the Gaia MultiLateral Agreement (MLA). The Gaia mission website is https://www.cosmos.esa.int/gaia. The Gaia archive website is https://archives.esac.esa.int/gaia.
Appendix A Wide Binary Table
Below, we provide a complete table of wide binaries with reported isochrone ages used in this work.
| Catalog | Gaia DR3 ID 1 | Gaia DR3 ID 2 | Age 1 | Age 2 | ||||
| [mas] | [mas] | [au] | [Gyr] | [Gyr] | ||||
| XR22aanot used in statistical analysis | 2684790761174460032 | 2684790658095155840 | 17590 | 0.000284 | ||||
| XR22 | 949512025266986368 | 949511995203757568 | 10407 | 0.000495 | ||||
| XR22 | 1331136194689123072 | 1331136194689123200 | 10348 | 0.000000 | ||||
| XR22 | 656616764883595264 | 656616769173736960 | 7201 | 0.000056 | ||||
| XR22 | 774012622101896832 | 774012622101896960 | 10004 | 0.000000 | ||||
| XR22 | 2098638466710584064 | 2098638466710583552 | 24694 | 0.001709 | ||||
| XR22 | 742452824453216000 | 742452824453215872 | 6987 | 0.000000 | ||||
| XR22 | 3283076097735358592 | 3283076097735358464 | 13466 | 0.000269 | ||||
| XR22 | 605026206128002432 | 605026304912470272 | 122214 | 0.000272 | ||||
| XR22 | 1271083343481863296 | 1271083343481863552 | 14891 | 0.000000 | ||||
| XR22 | 139325960240584448 | 139325960240584320 | 16145 | 0.000180 | ||||
| XR22 | 3712569838037565568 | 3712569838037565696 | 17012 | 0.000085 | ||||
| W23* | 2806661695148871808 | 2806661695148871680 | 4896 | 0.000033 | ||||
| W23* | 611480030145802496 | 611480030144420608 | 1143 | 0.000001 | ||||
| W23* | 3844873556811289600 | 3844873556811289472 | 4762 | 0.000001 | ||||
| W23* | 590400742973158528 | 590400742973157760 | 7474 | 0.000259 | ||||
| W23* | 433662989613654144 | 433662989613654528 | 3344 | 0.000179 | ||||
| W23* | 3340863061637513216 | 3340863061637513344 | 3651 | 0.000202 | ||||
| W23* | 2789349918929145472 | 2789349914633569408 | 2167 | 0.000002 | ||||
| W23* | 141380393419936512 | 141380393423544832 | 1310 | 0.002218 | ||||
| W23* | 3067850063855658368 | 3067850063855659648 | 6916 | 0.000162 | ||||
| W23* | 1244713171995779968 | 1244704307183280128 | 77120 | 0.050975 | ||||
| W23* | 384557288486804864 | 384557288488352640 | 1047 | 0.000066 | ||||
| W23* | 947313521704064128 | 947313521704064256 | 1918 | 0.000027 | ||||
| W23* | 2767876972235040640 | 2767876972235040768 | 2150 | 0.000000 | ||||
| W23* | 1402745600298908672 | 1402745600298908544 | 9034 | 0.000042 | ||||
| W23* | 3660638834410265600 | 3660638838705046528 | 3725 | 0.000725 | ||||
| W23* | 60121815498544384 | 60121815498544512 | 3258 | 0.000211 | ||||
| W23* | 718733231907234304 | 718733227611631488 | 5292 | 0.000003 | ||||
| W23* | 133883274603761792 | 133883274603761664 | 9563 | 0.002374 | ||||
| W23* | 3354320328049508096 | 3354320328049508480 | 8083 | 0.004061 | ||||
| W23* | 595918092322351744 | 595918092321353216 | 3939 | 0.000584 | ||||
| W23* | 339540052632252288 | 339539846473821824 | 36218 | 0.087791 | ||||
| W23 | 4419568802678498432 | 4419568802678498304 | 670 | 0.000000 | ||||
| W23 | 3441071589321145856 | 3441071585024494336 | 684 | 0.000024 | ||||
| W23 | 1021611714280235904 | 1021611714280235776 | 234 | 0.000000 | ||||
| W23 | 672661152082945920 | 672660774125826304 | 2487 | 0.000065 | ||||
| W23 | 47662012294520064 | 47661977934782336 | 3079 | 0.000015 | ||||
| W23 | 3674314048935804928 | 3674314048935804800 | 2329 | 0.000004 | ||||
| W23 | 4010176437034162688 | 4010176810697178752 | 299 | 0.000000 | ||||
| W23 | 866149428767077248 | 866149428768309632 | 193 | 0.000007 | ||||
| W23 | 3411752733808885888 | 3411751290699480576 | 17455 | 0.003017 | ||||
| W23 | 867326902642328704 | 867326902642328448 | 1855 | 0.000014 | ||||
| W23 | 1286001929561950336 | 1286001929563421696 | 190 | 0.000000 | ||||
| W23 | 57929561113486848 | 57875959921650688 | 109549 | 0.071794 | ||||
| W23 | 1535324570641254144 | 1535324570643080832 | 484 | 0.000009 | ||||
| W23 | 2513851749988035072 | 2513851749988035200 | 877 | 0.000001 | ||||
| W23 | 1157190706395007744 | 1157190706395008128 | 2600 | 0.000149 | ||||
| W23 | 3837150449699070208 | 3837150518418620032 | 4452 | 0.000450 | ||||
| W23 | 1209194857609837312 | 1209194857609837184 | 2903 | 0.000000 | ||||
| W23 | 949512025266986368 | 949511995203757568 | 10407 | 0.000495 | ||||
| W23 | 1495786755279354112 | 1495786755278552064 | 5012 | 0.000006 | ||||
| W23 | 173207411050004736 | 173207342330529152 | 9610 | 0.001188 | ||||
| W23 | 4573484350212176256 | 4573484350212177408 | 1560 | 0.000024 | ||||
| W23 | 382954849074369408 | 382955231327045888 | 525 | 0.000002 | ||||
| W23 | 664055515090902144 | 664055412011687424 | 1352 | 0.000009 | ||||
| W23 | 920594839393475840 | 920594835096999808 | 456 | 0.000001 | ||||
| W23 | 1527756464605583232 | 1527756464605583104 | 286 | 0.000000 | ||||
| W23 | 2863885056262025856 | 2863885056262025728 | 1441 | 0.000021 | ||||
| W23 | 826128369451071488 | 826128369452653952 | 4083 | 0.000000 | ||||
| W23 | 2682640906704081408 | 2682640902409329664 | 7585 | 0.015317 | ||||
| W23 | 1013649429029096192 | 1013649390375546368 | 3968 | 0.000053 | ||||
| W23 | 370866685094430720 | 370866685094430464 | 6442 | 0.000226 | ||||
| W23 | 2583631877487732480 | 2583631873193879680 | 2680 | 0.000025 | ||||
| W23 | 3306587057951531392 | 3306587057950933888 | 21862 | 0.001588 | ||||
| W23 | 3349903280603248512 | 3349903280603246336 | 9055 | 0.018810 | ||||
| W23 | 2725570135620008704 | 2725570131325603072 | 4081 | 0.005386 | ||||
| W23 | 3269518294730285312 | 3269518294730285184 | 3091 | 0.000000 | ||||
| W23 | 4017887239064594816 | 4017887239064738688 | 11093 | 0.002888 | ||||
| W23 | 92750166650167680 | 92750166650167552 | 3392 | 0.000000 | ||||
| W23 | 706100255620810752 | 706100186901333760 | 11915 | 0.001108 | ||||
| W23 | 605884679895137024 | 605884679895136896 | 23859 | 0.000003 | ||||
| W23 | 187983537364787712 | 187983541660709248 | 5518 | 0.000211 | ||||
| W23 | 78598691790870272 | 78598696086400000 | 2486 | 0.000000 | ||||
| W23 | 1223584200643469056 | 1223587151286176384 | 5750 | 0.000002 | ||||
| W23 | 142049141304409216 | 142049141304410368 | 24842 | 0.002834 | ||||
| W23 | 6898935568756589824 | 6898935564462716672 | 4918 | 0.000587 | ||||
| W23 | 655686651055845632 | 655686651055845504 | 11879 | 0.035775 | ||||
| W23 | 1290519719762528768 | 1290519715467767936 | 6748 | 0.000008 | ||||
| W23 | 4002494183650214784 | 4002494248074566656 | 19569 | 0.000000 | ||||
| W23 | 104567923783991808 | 104567885129642496 | 7589 | 0.001357 | ||||
| W23 | 182264019610189184 | 182264019610189440 | 10404 | 0.004093 | ||||
| W23 | 3142949044493143040 | 3142949044491195264 | 14226 | 0.085341 | ||||
| W23 | 1269273856578570624 | 1269273852283696128 | 2961 | 0.000000 | ||||
| W23 | 1538753401948398848 | 1538753397654113408 | 6955 | 0.000000 | ||||
| W23 | 3137368507588913920 | 3137368507588914176 | 24966 | 0.068164 | ||||
| W23 | 696086311735797376 | 696086316030823040 | 11006 | 0.010041 | ||||
| W23 | 376376028623902336 | 376376097343378688 | 14657 | 0.030784 | ||||
| N24* | 3819906254789418368 | 3819718586193769472 | 9591 | 0.000341 | ||||
| N24* | 6594232820498521344 | 6594232820498061312 | 5837 | 0.000007 | ||||
| N24* | 2153929092538747904 | 2153929053881531264 | 5228 | 0.000247 | ||||
| N24* | 3672659588877922176 | 3672659417079230592 | 8535 | 0.000006 | ||||
| N24* | 1331136194689123072 | 1331136194689123200 | 10348 | 0.000000 | ||||
| N24* | 5678381450362539648 | 5678384405300038912 | 36765 | 0.001999 | ||||
| N24* | 701934102984962048 | 701933931186270080 | 57160 | 0.046395 | ||||
| N24* | 6773067046729562112 | 6773066947947317760 | 61338 | 0.015659 | ||||
| N24* | 813941210770063232 | 813941215066621952 | 11299 | 0.000021 | ||||
| N24* | 6301399377571977472 | 6301399377571977728 | 13033 | 0.000034 | ||||
| N24* | 2233977559133346048 | 2233977176878402048 | 40292 | 0.002763 | ||||
| N24* | 3644460452815922048 | 3644460418456183424 | 12323 | 0.003498 | ||||
| N24* | 5292817739372153088 | 5292817739372153600 | 14115 | 0.010098 | ||||
| N24* | 4388903114145388544 | 4388903011066173696 | 16045 | 0.063508 | ||||
| N24* | 3159137325989430016 | 3159137321690419200 | 7364 | 0.000206 | ||||
| N24* | 4828407186261879296 | 4828407220621617536 | 30818 | 0.007727 | ||||
| N24 | 5335492530134122240 | 5335492740641614848 | 5179 | 0.000105 | ||||
| N24 | 4658108537614609536 | 4658107884688023040 | 9323 | 0.000447 | ||||
| N24 | 6068370532457290624 | 6068370528153600512 | 5557 | 0.000346 | ||||
| N24 | 3052542491179843840 | 3052542491179841408 | 11561 | 0.001048 | ||||
| N24 | 5950394785918358784 | 5950394751558620544 | 6295 | 0.000889 | ||||
| N24 | 1918228502880276608 | 1918228567301390720 | 13671 | 0.002197 | ||||
| N24 | 3002497154284865024 | 3002497154284866048 | 5890 | 0.000567 | ||||
| N24 | 3101720004155474304 | 3101719965495540736 | 7537 | 0.000986 | ||||
| N24 | 5352186758772274816 | 5352186797455149312 | 12842 | 0.000153 | ||||
| N24 | 4621120932002256256 | 4621026339642213376 | 27178 | 0.000786 | ||||
| N24 | 3265692990698073856 | 3265692986403196288 | 3004 | 0.000007 | ||||
| N24 | 2998185728674100608 | 2998185728674096768 | 19683 | 0.000606 | ||||
| N24 | 5405835543183097088 | 5405835646262319488 | 23314 | 0.000659 | ||||
| N24 | 6314750712909107072 | 6314750708613733504 | 5469 | 0.000762 | ||||
| N24 | 4224312580017776256 | 4224312580017776512 | 8461 | 0.010344 | ||||
| N24 | 6517920464353963392 | 6517920464353963008 | 22806 | 0.000000 | ||||
| N24 | 1736540283767059456 | 1736540322421822592 | 10020 | 0.000442 | ||||
| N24 | 5909557622697254272 | 5909745295588159744 | 11841 | 0.001145 | ||||
| N24 | 807831194591725568 | 807831194591725440 | 5535 | 0.000001 | ||||
| N24 | 5369418446746331136 | 5369418515465807232 | 15484 | 0.000000 | ||||
| N24 | 5906145425803385984 | 5906145421498047744 | 9001 | 0.000729 | ||||
| N24 | 1943687419745045504 | 1943687419745044992 | 26459 | 0.000763 | ||||
| N24 | 6342906246462442240 | 6342905494845063424 | 9882 | 0.013322 | ||||
| N24 | 3053054142045813120 | 3053054142045815552 | 10147 | 0.000046 | ||||
| N24 | 684440048352505856 | 684440048352505728 | 11376 | 0.000015 | ||||
| N24 | 5950052425488004480 | 5950051673901872512 | 14329 | 0.001540 | ||||
| N24 | 3102918712348299520 | 3102918815427513344 | 30382 | 0.007132 | ||||
| N24 | 998443694318500096 | 998443767332820736 | 18324 | 0.000058 | ||||
| N24 | 6771985303383361536 | 6771985337743100160 | 26065 | 0.005708 | ||||
| N24 | 2933769499933737216 | 2933769431214263680 | 15809 | 0.002395 | ||||
| N24 | 1005864435733114752 | 1005864440029460480 | 23504 | 0.020675 | ||||
| N24 | 3040821662865294336 | 3040821662865294208 | 6931 | 0.003592 | ||||
| N24 | 6238967320644157952 | 6238967247625278848 | 53579 | 0.058254 | ||||
| N24 | 1794586908506128640 | 1794586908506128384 | 7668 | 0.000122 | ||||
| N24 | 1407693574423019648 | 1407693574423023104 | 12584 | 0.001682 | ||||
| N24 | 5293455765353605120 | 5293455765353472896 | 15586 | 0.000445 | ||||
| N24 | 5547583822003400448 | 5547584032470737280 | 11756 | 0.001286 | ||||
| N24 | 5596320847444517760 | 5596321225401634944 | 17156 | 0.001131 | ||||
| N24 | 5361541850376235904 | 5361541850377466240 | 13862 | 0.000550 | ||||
| N24 | 5274203522910689408 | 5274203900867854720 | 15083 | 0.000607 | ||||
| N24 | 1938390144519877760 | 1938389972721184384 | 31193 | 0.001305 | ||||
| N24 | 5281398623923972224 | 5281398623923972992 | 36846 | 0.079655 | ||||
| N24 | 3107998112471916544 | 3107998112466660096 | 13719 | 0.002752 | ||||
| N24 | 1782755952936649472 | 1782755957232940288 | 13919 | 0.000585 | ||||
| N24 | 2104414854126478976 | 2104414819766738816 | 24657 | 0.006255 |
Note. — Columns: (1) Catalog name among Xiang & Rix 2022 (XR22), Wang et al. 2023 (W23), Nataf et al. 2024 (N24); () Gaia DR3 source ID of the primary and secondary; () Parallaxes of the primary and secondary in mas; (6) Projected separation of the binary in au; () Isochrone ages for the primary and secondary, (9) Chance alignment probability of the binary (; see El-Badry et al., 2021).